OpenStudy (anonymous):

How do you graph the inequality of this? -3y + 5 ≥ -22

OpenStudy (anonymous):

draw that line -3y + 5 =-22

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

I don't think It's that one.. But I don't know.

OpenStudy (anonymous):

-3y = -22-5 -3y = -27 y=27/3

OpenStudy (anonymous):

y= 9

OpenStudy (anonymous):



OpenStudy (anonymous):

then check origin 5 > -22

OpenStudy (anonymous):